Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
fafone, giri, lumaca, nardecchia, rocchi, menzione, polini, ruggi - 23:37 Thursday 18 February 2021 (50832) Print this report
NI CH alignment on CP - last round

Before starting the activity on the HWS-DET, we switched off the SLED for HWS-INJ. At 15.20 UTC, we disbabled the B5 drift control and checked the alignment of the optics. At 15.52 UTC, we moved the SR (SR_MIR_TY) by -600 µrad. No effect was observed on the HWS-DET after the SR misalignment and the illumination remained the same (see figure 1). A deeper analysis of the images will follow.

At 16.08 UTC, we switched on the CH with about 1,2 W on the CP and in around 10 mins, we observed the CH beam (figure 2), near where we left it in the last shift (50697).

At 16.22, we switch off the CH and switch on the RH at 16.49 UTC with P = 15W to see the center of the WI. After an hour, the center became clearly visible (figure 3).

We performed some iterations with the hot mirror to center the CH beam. Since the RH curvature (that we use as reference) was vanishing, we switched on the RH at 20:10 UTC, and the curvature (figure 4) appeared at 21.12 UTC. We switched off the RH and on again the CH.

When the CH curvature became visible on the HWS maps, we performed the last few operations to finalize the alignment (figure 5).

Images attached to this report
Search Help